# SafeEraser: Enhancing Safety in Multimodal Large Language Models through Multimodal Machine Unlearning

### v2

@@ Abstract
As Multimodal Large Language Models (MLLMs) develop, their potential security issues have become increasingly prominent. Machine Unlearning (MU), as an effective strategy for forgetting specific knowledge in training data, has been widely used in privacy protection. However, MU for safety in MLLM has yet to be fully explored. To address this issue, we propose SafeEraser, a safety unlearning benchmark for MLLMs, consisting of 3,000 images and 28.8K VQA pairs. We comprehensively evaluate unlearning methods from two perspectives: forget quality and model utility. Our findings show that existing MU methods struggle to maintain model performance while implementing the forget operation and often suffer from over-forgetting. Hence, we introduce Prompt Decouple (PD) Loss to alleviate over-forgetting through decouple prompt during unlearning process. To quantitatively measure over-forgetting mitigated by PD Loss, we propose a new metric called Safe Answer Refusal Rate (SARR). Experimental results demonstrate that combining PD Loss with existing unlearning methods can effectively prevent over-forgetting and achieve a decrease of 79.5% in the SARR metric of LLaVA-7B and LLaVA-13B, while maintaining forget quality and model utility. Our code and dataset will be released upon acceptance. Warning: This paper contains examples of harmful language and images, and reader discretion is recommended.
![[Uncaptioned image]](extracted/6230781/photo/logo.png)
@@ 1 Introduction
In recent years, Large Language Models (LLMs) (Dubey et al., 2024; Bai et al., 2023) and Multimodal Large Language Models (MLLMs) (Alayrac et al., 2022; Awadalla et al., 2023; Hurst et al., 2024) have achieved significant success in diverse applications Awais et al. (2025); Yan et al. (2024c, b). However, since model training involves millions of data, some of which may contain harmful knowledge, this could cause the model to generate harmful responses. Previous research (Luo et al., 2024; Liu et al., 2024c; Hu et al., 2024) has highlighted the severe security issues present in MLLMs. To confront the safety issue, numerous studies (Li et al., 2024b; Liu et al., 2024e; Łucki et al., 2024; Cooper et al., 2024) have revealed that Machine Unlearning (MU) could precisely forget harmful knowledge within LLMs.
@@ 2.1 MU for LLMs
The task of unlearning in LLMs has attracted significant attention in recent years (Barez et al., 2025). In previous studies, MU methods are typically divided into training-based methods and training-free methods. Training-based methods include gradient ascent (Bourtoule et al., 2020), gradient difference (Wang et al., 2023b; Yao et al., 2023), KL divergence (Yao et al., 2024), and preference optimization (Maini et al., 2024) and so on. Training-free methods include in-context unlearning (Pawelczyk et al., 2024) and corrupting prompt embeddings to achieve unlearning (Liu et al., 2024a). As MU methods for LLMs continue to evolve, constructing high-quality unlearning datasets and benchmarks has become increasingly important. Eldan and Russinovich (2023) propose a “Harry Potter” task for copyright, Maini et al. (2024) design an unlearning task with fictional authors, and Ma et al. (2024) introduce an unlearning benchmark for a fictional facial identity VQA dataset which aims to protect privacy. However, existing studies have not explored the application of MLLMs for forgetting harmful knowledge, a safety concern in MLLMs.
@@ 3.2.1 Dataset Statistics
Inspired by existing security benchmarks (Zhang et al., 2024; Gu et al., 2024), we define six major categories from different perspectives, with each category corresponding to 10 unique keywords and each keyword corresponding to 40 images. The dataset contains a total of 2,400 harmful images and 600 normal images, including 1,193 synthetic images and 1,807 real images. Additionally, the dataset contains a total of 28.8k QA pairs, which are divided into a forget set, a retain set, and a prompt decouple set, each containing 9,600 pairs.
@@ 3.2.2 Dataset Construction Pipeline
1) Concept-level: Concept-level data ensures the precise removal of harmful knowledge related to the keyword while retaining non-target knowledge associated with the keyword. We use GPT-4o to generate the keyword-related questions. The specific prompt used for GPT-4o be found in the Appendix G.1. We use vanilla model to generate the safe response as golden response.
@@ 3.3.1 Forget Quality
Attack Sucess Rate (ASR). When a MLLM processes a harmful text-image pair, the response is classified as either "safe" or "unsafe". A response is considered unsafe if it contains any harmful knowledge. We evaluate the model’s safety using the ASR, defined as , where is an indicator function that is 1 if the response contains harmful knowledge, and 0 otherwise and means the evaluation dataset.
Refusal Rate (RR). The previously defined ASR only measures the safety of the model’s responses. However, MLLLMs sometimes provide refusal responses to harmful queries. To assess the model’s ability to identify harmful queries and provide refusal responses, we introduce the RR, defined as , where is an indicator function that is 1 if the response contains refusal content, and 0 otherwise, means the evaluation dataset.
@@ 3.3.2 Model Utility
GPT-Eval. Inspired by LLM-as-a-Judge (Zheng et al., 2023), we use GPT-4o as an evaluator and introduce GPT-Eval, a metric that rates the correctness, helpfulness, and relevance of ’s responses on a scale from 0 to 1. we then expand the original scale to a range from 0 to 100. The detailed prompt for GPT-4o is in the Appendix G.2.
Specificity. Specificity evaluates the influence of unlearning methods on harmless knowledge. We employ several widely-used benchmarks to assess the Specificity of MLLMs. The detailed evaluation setting can be found in Appendix C.4.
@@ 5.2 Experiment Results
Attack Resistance. We use prompt-based and image-based attacks to evaluate the attack resistance of different MU methods. For prompt-based attacks, we employ Role-play Attack (RA), Ten-shot Prompting Attack (TPA), and Poetry Attack (PA). For image-based attacks, we use a method similar to Liu et al. (2024c). Detailed descriptions of these methods are provided in Appendix E. We use ASR to evaluate the performance of different MU methods on LLaVA-v1.5-7B. As shown in Table 3, we find that the methods combined with PD Loss achieve lower ASR, indicating that our approach offers better resistance to jailbreak attacks. This may be because PD Loss helps the model better distinguish between harmful and harmless knowledge, rather than simply "memorizing" harmful queries or images to achieve forgetting.
@@ B.2 Dataset Statistics
In the SafeEraser, we define six categories: Illegal Activity, Weapon, Violence, Hate Speech, Sex, and Privacy. Each category contains 10 relevant keywords, with a detailed list of categories and keywords provided in Figure 5. Within each category, there are 400 images, including both real and synthetic images. The specific number of real and synthetic images in each category can be found in Table 5.
@@ Appendix C Detailed Metric
C.2 Evaluation setting of Efficacy and Generality
C.3 Motivation of GPT-Eval
As shown by Wang et al. (2023a), traditional metrics like ROUGE often fail to capture semantic meaning. In our case, unlearning may reduce textual overlap, leading to lower ROUGE scores and high semantic similarity whichis acceptable.
@@ Appendix D Baseline MU methods
Appendix E Jailbreak Attack Methods
To evaluate the attack resistance of different MU methods, we design and implement several attack strategies from the perspectives of image-based and prompt-based attacks, aiming to access the model’s robustness under attack scenarios, as detailed below:
Image-based Attack: Following the approach of (Liu et al., 2024c), we generate an image with keywords on a white background and concatenate it with the original image, keeping the original query unchanged. This method aims to guide the model in recognizing and processing these additional texts through OCR.
Prompt-based Attacks: Prompt-based attacks embed harmful knowledge· into different forms of prompts. Following the approach of Doshi and Stickland (2024), Greshake et al. (2023) and Zou et al. (2023b), we use the following methods to access the performance of the unlearned model in prompt-based attack scenarios.
1) Poetry Attack (PA): We rephrase the query as poetry to increase linguistic complexity, evaluating whether the model can maintain unlearning performance with unconventional expressions. We used the following prompt: “Please keep the meaning of the query and express it in the form of a short poem: [Our Query],” converting the query into a poetic form.
2) Ten-shots Prompting Attack (TPA): We constructed 10 QA pairs from different domains as contextual input, covering topics like Biology, Geography, etc., so that the model could learn the format of normal responses, followed by inputting our harmful query, to evaluate if the model adapts to new examples.
3) Role-play Attack (RA): We create a role-playing scenario where characters introduce harmful questions in a dialogue, evaluating the model’s performance in a simulated conversation.

### v3

@@ Abstract
As Multimodal Large Language Models (MLLMs) develop, their potential security issues have become increasingly prominent. Machine Unlearning (MU), as an effective strategy for forgetting specific knowledge in training data, has been widely used in privacy protection. However, MU for safety in MLLM has yet to be fully explored. To address this issue, we propose SafeEraser, a safety unlearning benchmark for MLLMs, consisting of 3,000 images and 28.8K VQA pairs. We comprehensively evaluate unlearning methods from two perspectives: forget quality and model utility. Our findings show that existing MU methods struggle to maintain model performance while implementing the forget operation and often suffer from over-forgetting. Hence, we introduce Prompt Decouple (PD) Loss to alleviate over-forgetting through decouple prompt during unlearning process. To quantitatively measure over-forgetting mitigated by PD Loss, we propose a new metric called Safe Answer Refusal Rate (SARR). Experimental results demonstrate that combining PD Loss with existing unlearning methods can effectively prevent over-forgetting and achieve a decrease of 79.5% in the SARR metric of LLaVA-7B and LLaVA-13B, while maintaining forget quality and model utility. Our code and dataset can be found in https://github.com/yuu250/SafeEraser. Warning: This paper contains examples of harmful language and images, and reader discretion is recommended.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) (Dubey et al., 2024; Bai et al., 2023) and Multimodal Large Language Models (MLLMs) (Alayrac et al., 2022; Awadalla et al., 2023; Hurst et al., 2024) have achieved significant success in diverse applications Awais et al. (2025); Yan et al. (2024c, b). However, since model training involves millions of data, some of which may contain harmful knowledge, this could cause the model to generate harmful responses. Previous research (Luo et al., 2024; Liu et al., 2024c; Hu et al., 2024) has highlighted the severe security issues present in MLLMs. To confront the safety issue, numerous studies (Li et al., 2024b; Liu et al., 2024f; Łucki et al., 2024; Cooper et al., 2024) have revealed that Machine Unlearning (MU) could precisely forget harmful knowledge within LLMs.
@@ 2.1 MU for LLMs
The task of unlearning in LLMs has attracted significant attention in recent years (Barez et al., 2025). In previous studies, MU methods are typically divided into training-based methods and training-free methods. Training-based methods include gradient ascent (Bourtoule et al., 2020), gradient difference (Wang et al., 2023b; Yao et al., 2023), KL divergence (Yao et al., 2024), and preference optimization (Maini et al., 2024) and so on. Training-free methods include in-context unlearning (Pawelczyk et al., 2024) and corrupting prompt embeddings to achieve unlearning (Liu et al., 2024a). As MU methods for LLMs continue to evolve, constructing high-quality unlearning datasets and benchmarks has become increasingly important. Eldan and Russinovich (2023) propose a “Harry Potter” task for copyright, Maini et al. (2024) design an unlearning task with fictional authors, and Ma et al. (2024) Liu et al. (2024e) introduce an unlearning benchmark for a fictional facial identity VQA dataset which aims to protect privacy. However, existing studies have not explored the application of MLLMs for forgetting harmful knowledge, a safety concern in MLLMs.
@@ 3.2.1 Dataset Statistics
Inspired by existing security benchmarks (Zhang et al., 2024; Gu et al., 2024), we define six major categories, including Illegal Activity, Weapon, Violence, Hate Speech, Sex, Privacy, from different perspectives, with each category corresponding to 10 unique keywords and each keyword corresponding to 40 images. The detailed list of categories and keywords are provided in Figure 2. The dataset contains a total of 2,400 harmful images and 600 normal images, including 1,193 synthetic images and 1,807 real images. Additionally, the dataset contains a total of 28.8k QA pairs, which are divided into a forget set, a retain set, and a prompt decouple set, each containing 9,600 pairs.
@@ 3.2.2 Dataset Construction Pipeline
1) Concept-level: Concept-level data ensures the precise removal of harmful knowledge related to the keyword while retaining non-target knowledge associated with the keyword. We use GPT-4o to generate the keyword-related questions. The specific prompt used for GPT-4o can be found in the Appendix G.1. We use vanilla model to generate the safe response as golden response.
@@ 3.3.1 Forget Quality
Attack Success Rate (ASR). When a MLLM processes a harmful text-image pair, the response is classified as either "safe" or "unsafe". A response is considered unsafe if it contains any harmful knowledge. We evaluate the model’s safety using the ASR, defined as , where is an indicator function that is 1 if the response contains harmful knowledge, and 0 otherwise and means the evaluation dataset.
Refusal Rate (RR). The previously defined ASR only measures the safety of the model’s responses. However, MLLMs sometimes provide refusal responses to harmful queries. To assess the model’s ability to identify harmful queries and provide refusal responses, we introduce the RR, defined as , where is an indicator function that is 1 if the response contains refusal content, and 0 otherwise, means the evaluation dataset.
@@ 3.3.2 Model Utility
GPT-Eval. As shown by Wang et al. (2023a), traditional metrics like ROUGE often fail to capture semantic meaning. In our case, unlearning may reduce textual overlap, leading to lower ROUGE scores and high semantic similarity whichis acceptable. Inspired by LLM-as-a-Judge (Zheng et al., 2023), we use GPT-4o as an evaluator and introduce GPT-Eval, a metric that rates the correctness, helpfulness, and relevance of ’s responses on a scale from 0 to 1. We then expand the original scale to a range from 0 to 100. The detailed prompt for GPT-4o is in the Appendix G.2.
Specificity. Specificity evaluates the influence of unlearning methods on harmless knowledge. We employ several widely-used benchmarks to assess the Specificity of MLLMs, including GQA (Hudson and Manning, 2019), VisWiz (Gurari et al., 2018), SQA (Lu et al., 2022a), VQA (Lu et al., 2022b), POPE (Li et al., 2023b), Mm-Vet (Yu et al., 2023) and MMB (Liu et al., 2025b).
@@ 5.2 Experiment Results
Attack Resistance. We use prompt-based and image-based attacks to evaluate the attack resistance of different MU methods. For prompt-based attacks, we employ Role-play Attack (RA), Ten-shot Prompting Attack (TPA), and Poetry Attack (PA). For image-based attacks, we use a method similar to Liu et al. (2024c). The following are detailed explanations of these methods.
1) Image-based Attack: Following the approach of (Liu et al., 2024c), we generate an image with keywords on a white background and concatenate it with the original image, keeping the original query unchanged. This method aims to guide the model in recognizing and processing these additional texts through OCR.
2) Prompt-based Attacks: Prompt-based attacks embed harmful knowledge· into different forms of prompts. Following the approach of Doshi and Stickland (2024), Greshake et al. (2023) and Zou et al. (2023b), we use the following methods to access the performance of the unlearned model in prompt-based attack scenarios.
❶ Poetry Attack (PA): We rephrase the query as poetry to increase linguistic complexity, evaluating whether the model can maintain unlearning performance with unconventional expressions. We used the following prompt: “Please keep the meaning of the query and express it in the form of a short poem: [Our Query],” converting the query into a poetic form.
❶ Ten-shots Prompting Attack (TPA): We constructed 10 QA pairs from different domains as contextual input, covering topics like Biology, Geography, etc., so that the model could learn the format of normal responses, followed by inputting our harmful query, to evaluate if the model adapts to new examples.
❸ Role-play Attack (RA): We create a role-playing scenario where characters introduce harmful questions in a dialogue, evaluating the model’s performance in a simulated conversation.
Examples of attack methods are provided in the Appendix E. We use ASR to evaluate the performance of different MU methods on LLaVA-v1.5-7B. As shown in Table 3, we find that the methods combined with PD Loss achieve lower ASR, indicating that our approach offers better resistance to jailbreak attacks. This may be because PD Loss helps the model better distinguish between harmful and harmless knowledge, rather than simply "memorizing" harmful queries or images to achieve forgetting.
@@ 6 Conclusion
Acknowledgements
This work was supported by Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST (GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
@@ B.2 Dataset Statistics
In the SafeEraser, we define six categories: Illegal Activity, Weapon, Violence, Hate Speech, Sex, and Privacy. Each category contains 10 relevant keywords. Within each category, there are 400 images, including both real and synthetic images. The specific number of real and synthetic images in each category can be found in Table 5.
@@ Appendix D Baseline MU methods
Appendix E Example of Jailbreak Attack Methods